Menlo Company distributes a single product. The company’s sales and expenses for last month follow: Total Per Unit Sales $ 632,000 $ 40 Variable expenses 442,400 28 Contribution margin 189,600 $ 12 Fixed expenses 144,000 Net operating income $ 45,600 Required: 1. What is the monthly break-even point in unit sales and in dollar sales? 2. Without resorting to computations, what is the total contribution margin at the break-even point? 3-a. How many units would have to be sold each month to attain a target profit of $67,200? 3-b. Verify your answer by preparing a contribution format income statement at the target sales level. 4. Refer to the original data. Compute the company's margin of safety in both dollar and percentage terms. 5. What is the company’s CM ratio? If sales increase by $78,000 per month and there is no change in fixed expenses, by how much would you expect monthly net operating income to increase?
Answer #1

1) Break even Unit = Fixed cost/Contribution margin per unit = 144000/12 = 12000 Units

Break even Sales = 12000*40 = $480000

2) Total Contribution margin at break even = Fixed cost = $144000

3a) Required unit = (144000+67200)/12 = 17600 Units

3b) Contribution margin income statement

Sales (17600*40) 704000
Variable cost 492800
Contribution margin 211200
Fixed cost 144000
Net operating income 67200

4) Margin of safety (dollar) = 632000-480000 = 152000

Margin of safety (%) = 152000/632000 = 24.05%

5) CM ratio = 12/40 = 30%

Net operating income increase by = 78000*30% = 23400